# Designing Phase Sensitive Probes of Monopole Superconducting Order

### v1

@@ V Josephson junctions to probe $l_{z,\mathrm{loc}}$ of the monopole superconducting order
VI Josephson junctions to probe monopole superconducting order in the presence of spin-orbit interactions
@@
Lastly, in the presence of spin-orbit interactions at the junction interface or spin-orbit coupling in the bulk, the spin and orbital angular momentum are no longer conserved separately. The Josephson coupling is allowed when the pairing orders at two sides of the junction share the same value of total angular momentum along the common rotational symmetry axis following the junction geometry. For example, consider an SIS Josephson junction along $z$ axis between an $s$-wave superconductor with pairing $\Delta_{L,0}|\uparrow\downarrow-\downarrow\uparrow\rangle$ and a chiral $p$-wave triplet superconductor with zero total angular momentum pairing, for example $\Delta_{R,0}[(p_{x}-ip_{y}){|\uparrow\uparrow\rangle}+(p_{x}+ip_{y}){|\downarrow\downarrow\rangle}]$. In the presence of spin-orbit interactions, it shows nonzero first order Josephson coupling, similar to that in the Josephson junction of the same geometry between an $s$-wave superconductor and a monopole superconductor with pairing order $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$. Consequently, junction design (I) introduced in Sec. IV cannot distinguish $\Delta^{(q_{p}=-1,l_{z}=0)}_{\mathrm{MSC}}$ monopole pairing order from chiral $p$-wave pairing with $j_{z}=0$. Furthermore, junction design (II) introduced in Sec. V similarly cannot distinguish them, as the Majorana surface modes localized at the junction interface between chiral $p$-wave superconductors also give rise $4\pi$-periodic Josephson current phase relation. Hence, we must resort to a third junction design to discern the $\mathrm{\Delta}_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$ superconducting order in the presence of spin-orbit interaction at the barrier.
@@
As an example, we consider the effect of bare interface Rashba-type spin-orbit interaction. Generally, when spin-orbit interaction preserves time-reversal symmetry, the form factor between a spin singlet and a $q_{p}=-1$ monopole superconductor takes the following form,
@@
Now we consider a Josephson junction along $x$ direction between an $s$-wave superconductor and the $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$ monopole superconductor, as shown schematically in Fig. 7 (a). In the presence of Rashba-type spin-orbit interaction at the junction interface, $\mathbf{T}(\mathbf{k},\mathbf{k}^{\prime})=T_{\mathrm{SO}}\delta_{\mathbf{k}_{\parallel},\mathbf{k}^{\prime}_{\parallel}}\hat{\mathbf{x}}\times\hat{\mathbf{k}}$. For simplicity, we take the spin-independent tunneling amplitude as a constant, $T_{0}(\mathbf{k},\mathbf{k}^{\prime})=T_{0}$. Terms in the form factor in Eq. (41) that contribute to nonvanishing first order Josephson coupling are given by $\mathfrak{f}_{0}(\tilde{\mathbf{k}},\mathbf{k}^{\prime})=-T_{0}^{2}\Delta_{L,0}\Delta_{R,0}\sin^{2}\theta_{\tilde{k}}$ and $\mathfrak{f}_{2}(\tilde{\mathbf{k}},\mathbf{k}^{\prime})=-T_{\mathrm{SO}}^{2}|\hat{\mathbf{x}}\times\hat{\mathbf{k}}|^{2}\Delta_{L,0}\Delta_{R,0}\sin^{2}\theta_{\tilde{k}}\delta_{\mathbf{k}_{\parallel},\mathbf{k}^{\prime}_{\parallel}},$ with $\Delta_{L,0}$ and $\Delta_{R,0}$ being the pairing amplitudes of the $s$-wave and $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$ monopole superconductor respectively. Specifically, contributions from $\mathfrak{f}_{0}$ and $\mathfrak{f}_{2}$ are invariant under rotation $R_{x}$ and thus lead to nonzero first order Josephson current between the $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$ monopole superconductor and $s$-wave superconductor in a junction along the $x$ direction.
In contrast, a $(p_{x}-ip_{y}){|\uparrow\uparrow\rangle}+(p_{x}+ip_{y}){|\downarrow\downarrow\rangle}$ chiral $p$-wave superconductor has vanishing first order Josephson coupling with an $s$-wave superconductor in the same junction geometry, shown schematically in Fig. 7 (c), and with the same Rashba-type spin-orbit interaction at the interface. The form factor in Eq. (10) simplifies to $\mathfrak{F}_{\mathrm{sing}-\mathrm{trip}}\propto T_{0}T_{\mathrm{SO}}\hat{x}\cdot(\hat{\mathbf{k}}\times\mathbf{d}_{R}^{*}(\mathbf{k}))$. Here, the $d$-vector in consideration is $\mathbf{d}_{R}(\mathbf{k})\propto(k_{y},-k_{x},0)$. The form factor reduces to $\mathfrak{F}_{\mathrm{sing}-\mathrm{trip}}(\mathbf{k},\mathbf{k}^{\prime})\propto k_{x}k_{z}$, which is odd under mirror $M_{z}$ reflection, $k_{z}\mapsto-k_{z}$. Therefore, the net first order Josephson current of this system is zero. As such, the presence of first order Josephson current in this additional junction design can be used to distinguish $(p_{x}-ip_{y}){|\uparrow\uparrow\rangle}+(p_{x}+ip_{y}){|\downarrow\downarrow\rangle}$ chiral superconductor from the $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$ monopole superconductor. For other monopole superconducting pairing orders in different $q_{p}$ or $l_{z}$ sectors, one can design similar junctions to distinguish monopole superconducting order from chiral pairing orders in the presence of bare spin-orbit interactions at the interface.
@@
We demonstrate the analytic results above using numerical calculation of the Josephson current based on tight binding models. The system is analogous to that introduced in Sec. IV.1.2, only now the junction is oriented along $x$ direction and includes spin-dependent tunneling at the interface. For given conserved transverse momenta $k_{y}$ and $k_{z}$, we model the junction barrier as
@@
Above, $\hat{t}(k_{y},k_{z})=t_{0}\sigma_{0}+\mathbf{t}(k_{y},k_{z})\cdot\bm{\sigma}$, in which $t_{0}$ is the spin-independent hopping amplitude, and $\mathbf{t}(k_{y},k_{z})=t_{\mathrm{SO}}(\sin k_{y}\hat{z}-\sin k_{z}\hat{y})$ is the Rashba-type spin-orbit coupling at the interface, with $t_{\mathrm{SO}}$ being the spin-dependent hopping amplitude.
@@
The Josephson current for the junction along $x$ direction between a $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$ monopole superconductor and $s$-wave superconductor is shown in Fig. 7 (b). In agreement with the above arguments using the form factor $\mathfrak{F}_{\mathrm{sing}-\mathrm{MSC}}^{(q_{p}=-1)}$, there is an overall $2\pi$-periodic Josephson current corresponding to nonvanishing first order Josephson coupling. In contrast, for a junction between an $s$-wave superconductor and $(p_{x}-ip_{y}){|\uparrow\uparrow\rangle}+(p_{x}+ip_{y})\left|\downarrow\downarrow\right\rangle$ chiral superconductor, there is no first order Josephson coupling, as shown in Fig. 7 (d). Rather, the leading order Josephson coupling between the $s$-wave and chiral $p$-wave superconductor is a second order process, corresponding to $\pi$-periodic Josephson current. It is to be noted that because we have taken the spin-independent and spin-dependent tunneling amplitudes to both be of order $1$ in the tight-binding model, the magnitude of the Josephson critical current for the two junctions is similar; however, this does not change the qualitative aspects, specifically the periodicity. The numerical results are in agreement with the above microscopic analysis and demonstrate that this additional junction can be used to distinguish $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$ pairing order from other chiral $p$-wave superconductors in the presence of spin-orbit interaction at the interface. Similar results hold for a chiral $(p_{x}+ip_{y})|\downarrow\downarrow\rangle$ superconductor, $(p_{x}-ip_{y})\left|\uparrow\uparrow\right\rangle$ superconductor, or linear superposition of the two. In other words, the monopole superconducting order cannot be described as a simple linear superposition of spin-polarized chiral superconducting pairing orders globally but rather exists in its own distinct topological class.
VII Conclusion
@@
We have proposed using a set of phase-sensitive Josephson junctions to probe monopole superconducting orders and to distinguish them from other known pairing orders with spherical harmonic symmetry. We develop the designs based on two approaches. Firstly, we employed the linear response approach to derive the first order Josephson current between two uniform superconductors. In addition to accounting for the microscopic details of the tunneling processes, including spin-orbit interactions at the junction interface and spin-orbit coupling in the bulk superconductors, we emphasize that the symmetry of superconducting pairing orders and the junction geometry is encoded in a form factor $\mathfrak{F}(\mathbf{k},\mathbf{k}^{\prime})$, which is nonsingular and determines the symmetry selection rules of the first order Josephson coupling. Complementing the first approach, we have also proposed symmetry and topological principles in the design of phase-sensitive probes of monopole superconducting order. We noted that the $\mathrm{U}(1)$ topological obstruction of the pairing phase can be exhibited as a nonzero shift between the global and local angular momentum of the monopole superconducting order. Then, following symmetry principles, we proposed two classes of Josephson junctions of monopole superconductors to probe the global and local angular momentum of the superconducting order.
@@
We considered the first class of Josephson junctions aligned along the rotational axis ($z$ axis) between a monopole superconductor and other superconductors with known pairing symmetry. As the first order $2\pi$-periodic Josephson current is allowed only when superconducting orders at both sides of the junction transform under rotation $R_{z}$ in the same way, we can use this to probe $l_{z,\mathrm{glob}}$, the global angular momentum of the monopole pairing order. We have studied Josephson currents in this class of junctions between monopole superconductors with different pair monopole charges and different $l_{z,\mathrm{glob}}$ and superconductors with standard $s$- and $d$-wave spherical harmonic symmetry, using both linear response and tight-binding calculations. The second class of junctions, designed to extract the local angular momentum $l_{z,\mathrm{loc}}$ of the pairing order, consists of two identical monopole superconductors in a junction aligned perpendicular to the rotational $z$ axis. The Josephson current can be viewed as a superposition of contributions from different values of conserved momentum $k_{z}$ local in momentum space. By analyzing the existence of $4\pi$-periodic Josephson current at different orientations of the monopole superconductor based on tight-binding models, we have identified the local chiral symmetry of monopole pairing order characterized by angular momentum $l_{z,\mathrm{loc}}$. The shift between $l_{z,\mathrm{glob}}$ and $l_{z,\mathrm{loc}}$ indicates the nonzero pair monopole charge of the exotic superconducting order. The underlying symmetry and topological principles of this work provide useful guidance for the ongoing experimental efforts toward realizing and identifying exotic monopole superconducting pairing orders in doped topological semimetal materials.
Acknowledgements
@@
Appendix A Linear response theory of Josephson coupling in presence of spin-orbit interactions
We review the derivation of the first order Josephson current in an SIS junction via linear response theory, accounting for a general form of spin-orbit interaction in the bulk superconductors or at the junction interface. We consider the Hamiltonian of the Josephson junction presented in Eq. (1), in which two superconductors are connected by a weak link modelled by the tunneling Hamiltonian in Eq. (3).
A.1 Evaluating first order Josephson coupling
The total electric tunneling current through the junction reads $I(t)=|e|\langle\dot{N}_{L}(t)\rangle$, in which $N_{L}=\sum_{\sigma,\mathbf{k}}c^{\dagger}_{L,\mathbf{k},\sigma}c_{L,\mathbf{k},\sigma}$ is the number of particles at the left side of the barrier. We evaluate using linear response theory, in which $H_{\mathrm{link}}$ in Eq. (3) is treated perturbatively in the thermal weight expansion to the first order. The first order current is given by $I(t)={-i(e/\hbar)\int_{-\infty}^{+\infty}\mathrm{d}t^{\prime}\Theta(t-t^{\prime})\langle[\dot{N}_{L}(t),H_{\mathrm{link}}(t^{\prime})]\rangle}$, where $\Theta(t-t^{\prime})$ is the Heaviside function, and $\dot{N}_{L}(t)$ and $H_{\mathrm{link}}(t^{\prime})$ are in the interaction picture. It follows that the current is a superposition of the normal current from quasiparticle tunneling and the Josephson supercurrent from the tunneling of Cooper pairs [44]. The Josephson tunneling current is given by the imaginary part of the current-current correlation function,

### v2

@@ VI Long Josephson junctions
VII Josephson junctions to probe monopole superconducting order in the presence of spin-orbit interactions
@@
Lastly, in the presence of spin-orbit interactions at the junction interface or spin-orbit coupling in the bulk, the spin and orbital angular momentum are no longer conserved separately. The Josephson coupling is allowed when the pairing orders at two sides of the junction share the same value of total angular momentum along the common rotational symmetry axis following the junction geometry. For example, consider an SIS Josephson junction along $z$ axis between an $s$-wave superconductor with pairing $\Delta_{L,0}|\uparrow\downarrow-\downarrow\uparrow\rangle$ and a chiral $p$-wave triplet superconductor with zero total angular momentum pairing, for example $\Delta_{R,0}[(p_{x}-ip_{y}){|\uparrow\uparrow\rangle}+(p_{x}+ip_{y}){|\downarrow\downarrow\rangle}]$. In the presence of spin-orbit interactions, it shows nonzero first-order Josephson coupling, similar to that in the Josephson junction of the same geometry between an $s$-wave superconductor and a monopole superconductor with pairing order $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$. Consequently, junction design (I) introduced in Sec. IV cannot distinguish $\Delta^{(q_{p}=-1,l_{z}=0)}_{\mathrm{MSC}}$ monopole pairing order from chiral $p$-wave pairing with $j_{z}=0$. Furthermore, junction design (II) introduced in Sec. V similarly cannot distinguish them, as the Majorana surface modes localized at the junction interface between chiral $p$-wave superconductors also give rise $4\pi$-periodic Josephson current phase relation. Hence, we must resort to a third junction design to discern the $\mathrm{\Delta}_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$ superconducting order in the presence of spin-orbit interaction at the barrier.
@@
As an example, we consider the effect of bare interface Rashba-type spin-orbit interaction. Generally, when spin-orbit interaction preserves time-reversal symmetry, the form factor between a spin singlet and a $q_{p}=-1$ monopole superconductor takes the following form:
@@
Now we consider a Josephson junction along $x$ direction between an $s$-wave superconductor and the $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$ monopole superconductor, as shown schematically in Fig. 8(a). In the presence of Rashba-type spin-orbit interaction at the junction interface, $\mathbf{T}(\mathbf{k},\mathbf{k}^{\prime})=T_{\mathrm{SO}}\delta_{\mathbf{k}_{\parallel},\mathbf{k}^{\prime}_{\parallel}}\hat{\mathbf{x}}\times\hat{\mathbf{k}}$. For simplicity, we take the spin-independent tunneling amplitude as a constant, $T_{0}(\mathbf{k},\mathbf{k}^{\prime})=T_{0}$. Terms in the form factor in Eq. (42) that contribute to nonvanishing first-order Josephson coupling are given by $\mathfrak{f}_{0}(\tilde{\mathbf{k}},\mathbf{k}^{\prime})=-T_{0}^{2}\Delta_{L,0}\Delta_{R,0}\sin^{2}\theta_{\tilde{k}}$ and $\mathfrak{f}_{2}(\tilde{\mathbf{k}},\mathbf{k}^{\prime})=-T_{\mathrm{SO}}^{2}|\hat{\mathbf{x}}\times\hat{\mathbf{k}}|^{2}\Delta_{L,0}\Delta_{R,0}\sin^{2}\theta_{\tilde{k}}\delta_{\mathbf{k}_{\parallel},\mathbf{k}^{\prime}_{\parallel}},$ with $\Delta_{L,0}$ and $\Delta_{R,0}$ being the pairing amplitudes of the $s$-wave and $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$ monopole superconductor respectively. Specifically, contributions from $\mathfrak{f}_{0}$ and $\mathfrak{f}_{2}$ are invariant under rotation $R_{x}$ and thus lead to nonzero first-order Josephson current between the $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$ monopole superconductor and $s$-wave superconductor in a junction along the $x$ direction.
In contrast, a $(p_{x}-ip_{y}){|\uparrow\uparrow\rangle}+(p_{x}+ip_{y}){|\downarrow\downarrow\rangle}$ chiral $p$-wave superconductor has vanishing first-order Josephson coupling with an $s$-wave superconductor in the same junction geometry, shown schematically in Fig. 8(c), and with the same Rashba-type spin-orbit interaction at the interface. The form factor in Eq. (10) simplifies to $\mathfrak{F}_{\mathrm{sing}-\mathrm{trip}}\propto T_{0}T_{\mathrm{SO}}\hat{x}\cdot(\hat{\mathbf{k}}\times\mathbf{d}_{R}^{*}(\mathbf{k}))$. Here, the $d$-vector in consideration is $\mathbf{d}_{R}(\mathbf{k})\propto(k_{y},-k_{x},0)$. The form factor reduces to $\mathfrak{F}_{\mathrm{sing}-\mathrm{trip}}(\mathbf{k},\mathbf{k}^{\prime})\propto k_{x}k_{z}$, which is odd under mirror $M_{z}$ reflection, $k_{z}\mapsto-k_{z}$. Therefore, the net first-order Josephson current of this system is zero. As such, the presence of first-order Josephson current in this additional junction design can be used to distinguish $(p_{x}-ip_{y}){|\uparrow\uparrow\rangle}+(p_{x}+ip_{y}){|\downarrow\downarrow\rangle}$ chiral superconductor from the $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$ monopole superconductor. For other monopole superconducting pairing orders in different $q_{p}$ or $l_{z}$ sectors, one can design similar junctions to distinguish monopole superconducting order from chiral pairing orders in the presence of bare spin-orbit interactions at the interface.
@@
We demonstrate the analytic results above using numerical calculation of the Josephson current based on tight-binding models. The system is analogous to that introduced in Sec. IV.1.2, only now the junction is oriented along $x$ direction and includes spin-dependent tunneling at the interface. For given conserved transverse momenta $k_{y}$ and $k_{z}$, we model the junction barrier as
@@
Here, $\hat{t}(k_{y},k_{z})=t_{0}\sigma_{0}+\mathbf{t}(k_{y},k_{z})\cdot\bm{\sigma}$, where $t_{0}$ is the spin-independent hopping amplitude, and $\mathbf{t}(k_{y},k_{z})=t_{\mathrm{SO}}(\sin k_{y}\hat{z}-\sin k_{z}\hat{y})$ is the Rashba-type spin-orbit coupling at the interface, with $t_{\mathrm{SO}}$ being the spin-dependent hopping amplitude.
@@
The Josephson current for the junction along $x$ direction between a $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$ monopole superconductor and $s$-wave superconductor is shown in Fig. 8(b). In agreement with the above arguments using the form factor $\mathfrak{F}_{\mathrm{sing}-\mathrm{MSC}}^{(q_{p}=-1)}$, there is an overall $2\pi$-periodic Josephson current corresponding to nonvanishing first-order Josephson coupling. In contrast, for a junction between an $s$-wave superconductor and $(p_{x}-ip_{y}){|\uparrow\uparrow\rangle}+(p_{x}+ip_{y})\left|\downarrow\downarrow\right\rangle$ chiral superconductor, there is no first-order Josephson coupling, as shown in Fig. 8(d). Rather, the leading order Josephson coupling between the $s$-wave and chiral $p$-wave superconductor is a second-order process, corresponding to $\pi$-periodic Josephson current. It is to be noted that because we have taken the spin-independent and spin-dependent tunneling amplitudes to both be of order $1$ in the tight-binding model, the magnitude of the Josephson critical current for the two junctions is similar; however, this does not change the qualitative aspects, specifically the periodicity. The numerical results are in agreement with the above microscopic analysis and demonstrate that this additional junction can be used to distinguish $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$ pairing order from other chiral $p$-wave superconductors in the presence of spin-orbit interaction at the interface. Similar results hold for a chiral $(p_{x}+ip_{y})|\downarrow\downarrow\rangle$ superconductor, $(p_{x}-ip_{y})\left|\uparrow\uparrow\right\rangle$ superconductor, or linear superposition of the two. In other words, the monopole superconducting order cannot be described as a simple linear superposition of spin-polarized chiral superconducting pairing orders globally but rather exists in its own distinct topological class.
VIII Conclusion
@@
We have proposed using a set of phase-sensitive Josephson junctions to probe monopole superconducting orders and to distinguish them from other known pairing orders with spherical harmonic symmetry. We develop the designs based on two approaches. Firstly, we employed the linear response approach to derive the first-order Josephson current between two uniform superconductors. In addition to accounting for the microscopic details of the tunneling processes, including spin-orbit interactions at the junction interface and spin-orbit coupling in the bulk superconductors, we emphasize that the symmetry of superconducting pairing orders and the junction geometry is encoded in a form factor $\mathfrak{F}(\mathbf{k},\mathbf{k}^{\prime})$, which is nonsingular and determines the symmetry selection rules of the first-order Josephson coupling. Complementing the first approach, we have also proposed symmetry and topological principles in the design of phase-sensitive probes of monopole superconducting order. We noted that the $\mathrm{U}(1)$ topological obstruction of the pairing phase can be exhibited as a nonzero shift between the global and local angular momentum of the monopole superconducting order. Then, following symmetry principles, we proposed two classes of Josephson junctions of monopole superconductors to probe the global and local angular momentum of the superconducting order.
@@
We considered the first class of Josephson junctions aligned along the rotational axis ($z$ axis) between a monopole superconductor and other superconductors with known pairing symmetry. As the first-order $2\pi$-periodic Josephson current is allowed only when superconducting orders at both sides of the junction transform under rotation $R_{z}$ in the same way, we can use this to probe $l_{z,\mathrm{glob}}$, the global angular momentum of the monopole pairing order. We have studied Josephson currents in this class of junctions between monopole superconductors with different pair monopole charges and different $l_{z,\mathrm{glob}}$ and superconductors with standard $s$- and $d$-wave spherical harmonic symmetry, using both linear response and tight-binding calculations. The second class of junctions, designed to extract the local angular momentum $l_{z,\mathrm{loc}}$ of the pairing order, consists of two identical monopole superconductors in a junction aligned perpendicular to the rotational $z$ axis. The Josephson current can be viewed as a superposition of contributions from different values of conserved momentum $k_{z}$ local in momentum space. By analyzing the existence of $4\pi$-periodic Josephson current at different orientations of the monopole superconductor based on tight-binding models, we have identified the local chiral symmetry of monopole pairing order characterized by angular momentum $l_{z,\mathrm{loc}}$. The shift between $l_{z,\mathrm{glob}}$ and $l_{z,\mathrm{loc}}$ indicates the nonzero pair monopole charge of the exotic superconducting order. The underlying symmetry and topological principles of this work provide useful guidance for the ongoing experimental efforts toward realizing and identifying exotic monopole superconducting pairing orders in doped topological semimetal materials.
Acknowledgments
@@
Appendix A Linear response theory of Josephson coupling in the presence of spin-orbit interactions
We review the derivation of the first-order Josephson current in an SIS junction via linear response theory, accounting for a general form of spin-orbit interaction in the bulk superconductors or at the junction interface. We consider the Hamiltonian of the Josephson junction presented in Eq. (1), where two superconductors are connected by a weak link modeled by the tunneling Hamiltonian in Eq. (3).
A.1 First-order Josephson current and form factor in the presence of spin-dependent tunneling
The total electric tunneling current through the junction is $I(t)=|e|\langle\dot{N}_{L}(t)\rangle$, where $N_{L}=\sum_{\sigma,\mathbf{k}}c^{\dagger}_{L,\mathbf{k},\sigma}c_{L,\mathbf{k},\sigma}$ is the number of particles at the left side of the barrier. We evaluate using linear response theory, where $H_{\mathrm{link}}$ in Eq. (3) is treated perturbatively in the thermal weight expansion to the first-order. the first-order current is given by $I(t)={-i(e/\hbar)\int_{-\infty}^{+\infty}\mathrm{d}t^{\prime}\Theta(t-t^{\prime})\langle[\dot{N}_{L}(t),H_{\mathrm{link}}(t^{\prime})]\rangle}$, where $\Theta(t-t^{\prime})$ is the Heaviside function, and $\dot{N}_{L}(t)$ and $H_{\mathrm{link}}(t^{\prime})$ are in the interaction picture. It follows that the current is a superposition of the normal current from quasiparticle tunneling and the Josephson supercurrent from the tunneling of Cooper pairs Mahan (2000). The Josephson tunneling current is given by the imaginary part of the current-current correlation function,